When I am as old as sue is now, sue will be 4 years older than I am now. How much older is sue than I? Please if you can explain how you got it because I'm having a hard time with it and I would like to understand. Thank you!
Found 3 solutions by ankor@dixie-net.com, ikleyn, greenestamps:
Answer by ankor@dixie-net.com(22740)

When I am as old as sue is now, sue will be 4 years older than I am now.
How much older is sue than I?
:
let y = your age now
let s = Sue's age now
let d = s - y, the difference between your age and Sue's age
"When I am as old as sue is now, "
y + d = s
sue will be 4 years older than I am now.
s = y + 4
s = s therefore
y + d = y + 4
d = 4
Replace d with (s-y)
s - y = 4
Obviously Sue is 4 yrs older than you

Answer by ikleyn(52803)

.

When I am as old as sue is now, sue will be 4 years older than I am now.
How much older is sue than I?
~~~~~~~~~~~~~~~~~~~~~~~~~


            The solution by the other tutor is  TOTALLY WRONG,  so I came to fix it.

            Below find my correct solution.


Let y = your age now.

Let s = Sue's age now.

Let d = s - y,     (*)  the difference between Sue's age and your age.


"When I am as old as sue is now, "     (- It will happen in d years from now, and Sue will be (s + d) years old then)


sue will be 4 years older than I am now.
s + d = y + 4     (**)           <<<---===  The other tutor missed "+ d" in the left side of this equation,
                                            therefore all his further conclusions are wrong


From (**),  s - y = 4 - d = (substitute here d= s - y   from (*)) = 4 - (s-y).


Thus you have  s- y = 4 - (s-y).


Collect the terms  "(s-y)" on the left side:


2*(s-y) = 4,   


hence,  s-y = 2,


which means that Sue is 2 years older than you.     ANSWER

Solved.


Answer by greenestamps(13200)



For many students, a formal algebraic solution like the one by tutor @ikleyn is probably the best way to solve the problem.

For some students who are good with logical reasoning, you can solve the problem in basically the same way, but without the formal algebra.

In some number of years x, I will be as old as Sue is now.
After that number of years, Sue will be x years older than SHE is now.
If AT THAT TIME Sue will be 4 years older than I am NOW, then those 4 years will be equal to 2 times x.

So x is 2; since that is the number of years it will be before I am as old as Sue, Sue is 2 years older than I am.